The ages of Sam, Tonton, and Ramon are consecutive integers.Sam is younger that Ramon but older that Tonton.Sam's Mother is Thrice as old as Sam and 14 years older than twice as old as Ramon.Find the ages of each of them
Answer by ankor@dixie-net.com(22740)

The ages of Sam, Tonton, and Ramon are consecutive integers.
Sam is younger than Ramon but older that Tonton.
Sam's Mother is Thrice as old as Sam and 14 years older than twice as old as Ramon.
Find the ages of each of them
:
Let the three ages be: x, (x+1), and (x+2)
let m = Sam's Mom
:
Determine which is which
"Sam is younger than Ramon but older that Tonton." therefore:
x = Ton's age
(x+1) = Sam's age
(x+2) = Ram's age
:
"Sam's Mother is Thrice as old as Sam":
m = 3(x+1)
"and 14 years older than twice as old as Ramon.
m = 2(x+2) + 14
m = m therefore
3(x+1) = 2(x+2) + 14
3x + 3 = 2x + 4 + 14
3x - 2x = 4 + 14 - 3
x = 15 yrs is Ton's age
then obviously
16 = Sam's age
17 = Ram's age
then
m = 3(16)
m = 48 yrs is Mom's age
:
;
See if the that checks out
Sam's Mother is 14 years older than twice as old as Ramon.
m = 2(17) + 14
m = 48